Bobby just received an award for employee of the year and want to frame the plaque in such a way as it creates a border of uniform width around the rectangular plaque. He only has enough material to create a total of 16 square inches in area for the border. If the plaque 6 inches by 9 inches, how wide should the border be?
thank you!

Answer by Boreal(15235)

The area of the plaque is 54 in^2
The area of the total is 70 in^2, adding the border.
The dimensions of the border are (6+2x)(9+2x), x being the width.
This has an area of 54+30x+4x^2=70
so 4x^2+30x-16=0
or 2x^2+15x-8=0
(2x-1)(x+8)=0
positive root only is x=(1/2) inch
The plaque and border are 7*10 inches